Deduce the ratio of the mass of oxygen per gram of sulfur in the compound sulfur dioxide
1
Expert's answer
2020-03-23T10:13:18-0400

Sulfur dioxide"-SO_2"

Mass of sulfur in the compound "=32 \ g"

Mass of oxygen in the compound "=2\times 16=32\ g"

So,mass of oxygen per gram of sulfur will be "\ \frac{32}{32}\times 1=1 \ g"

The ratio will be "1:1" .